The diagonal of a rectangle exceeds the length by 1. The width of the rectangle is 1 less than one-third of the length. Find the length of the diagonal. Please show all steps.
Thanks :)

Answer by richwmiller(17219)

L,w,d
d=L+1
w=l/3-1
L^2+w^2=d^2
In the last equation
substitute L+1 for d
substitute (L/3)-1
solve for L
d=25
l=24
w=7
Have fun!
:)